# Eye-gaze Guided Multi-modal Alignment Framework for Radiology

### v2

@@ Abstract
In multi-modal frameworks, the alignment of cross-modal features presents a significant challenge. The predominant approach in multi-modal pre-training emphasizes either global or local alignment between modalities, utilizing extensive datasets. This bottom-up driven method often suffers from a lack of interpretability, a critical concern in radiology. Previous studies have integrated high-level labels in medical images or text, but these still rely on manual annotation, a costly and labor-intensive process. Our work introduces a novel approach by using eye-gaze data, collected synchronously by radiologists during diagnostic evaluations. This data, indicating radiologists’ focus areas, naturally links chest X-rays to diagnostic texts. We propose the Eye-gaze Guided Multi-modal Alignment (EGMA) framework to harness eye-gaze data for better alignment of image and text features, aiming to reduce reliance on manual annotations and thus cut training costs. Our model demonstrates robust performance, outperforming other state-of-the-art methods in zero-shot classification and retrieval tasks. The incorporation of easily-obtained eye-gaze data during routine radiological diagnoses signifies a step towards minimizing manual annotation dependency. Additionally, we explore the impact of varying amounts of eye-gaze data on model performance, highlighting the feasibility and utility of integrating this auxiliary data into multi-modal pre-training.
@@ I Introduction
With the development of multi-modal learning, models in the pre-training phase can now utilize large amounts of paired multi-modal data, such as image-text pairs, audio-text pairs, etc., to optimize the multi-modal feature extraction and alignment capabilities. With the emergence of the CLIP [1] model, contrastive learning has become the prominent framework of multi-modal learning. The advantage of this framework lies in its simplicity of structure and it does not require sample-level annotations. However, the main drawback is its heavy reliance on the scale of training data. When trained on smaller-scale data, the model cannot accurately learn the alignment relationship between multi-modal features. Subsequent works have sought to optimize this framework. For instance, GLIP [2] and RegionCLIP [3] utilized pre-predicted annotation information to perform fine-grained region-level pre-training. They used detection networks firstly to predict image regions relevant to the text prompt, and then trained the model to align these image regions with their corresponding text descriptions. However, these models heavily rely on the performance of the ROI detector and have high computational complexity. FILIP [4] proposed a refined multi-modal alignment method using image patches and text tokens, choosing to perform this operation after the encoder. This significantly reduces the burden on training the model and enhances the flexibility of the architecture.
In the medical domain, there are also datasets available for training multi-modal models, such as MIMIC-CXR [5] containing 377,110 chest images and 227,827 corresponding diagnostic reports, and CheXpert [6] containing 224,316 chest images with semantic labels. While these datasets are smaller in scale compared to natural image datasets, models pretrained on them are better suited for medical tasks like classification, retrieval, and VQA. To address the scarcity of medical data, studies [7, 8] have introduced self-supervised training into the CLIP framework to further enhance encoder performance. Additionally, weak labels between images and texts have been incorporated during pretraining to aid multi-modal alignment [9]. Some studies [10, 11] utilized fine-grained alignment between chest image patches and text tokens for pretraining, similar to approaches [4] in natural image domains. However, unlike natural images, the relationship between medical images and diagnostic texts is often more complex and challenging to learn, and in some domains, the available datasets for training remain small in scale. Consequently, models trained in the pretraining phase are prone to over-fitting, learning non-essential shortcut [12] features, thereby diminishing the model’s generalization and interpretability.
@@
In this study, we fully explore the auxiliary role of eye-gaze data from radiologists. Eye-gaze data can intuitively reflect the image regions radiologists focus on, providing insights into their cognitive behavior during diagnosis. When collecting eye-gaze data simultaneously with radiologists’ observation of images and verbalization of diagnostic results, it serves as a bridge for aligning images and text, as radiologists tend to examine relevant image areas when providing diagnoses (as illustrated in Fig. 1). Therefore, we believe this data can serve as an auxiliary role in aligning textual and visual features. For the data collection, study [13] developed an eye-tracking system that allows radiologists to interact freely with images, departing from the fixed-image acquisition mode, thereby offering a feasible solution for data collection in routine work settings. Moreover, such a system significantly reduces the time radiologists spend annotating data, laying the groundwork for future data expansion. And there already exists a publicly available medical multi-modal dataset called MIMIC-EYE [14]. Thus, considering the utilization of eye-tracking data to assist in multi-modal model training, we propose the Eye-gaze Guided Multi-modal Alignment framework (EGMA). Our model first segments the transcribed text into individual sentences and obtains radiologists’ attention heatmaps. Subsequently, we obtain encoded features of image patches and sentences through image and text encoders, generating instance-level similarity matrix. Then, we compute the loss between this matrix and the attention heatmaps, integrating refined feature representations for subsequent contrastive loss. To further leverage the assisting role of eye-gaze data in aligning images and texts, we combine the eye-gaze heatmaps with the similarity matrix derived from model, serving as weights to calculate cross-modal mapping loss. Experimental results on zero-shot classification and retrieval tasks reveal that our framework surpasses other leading methods in performance across diverse datasets and under multiple dataset size scenarios. Specifically, the EGMA framework yielded a remarkable 3.9% improvement in image-to-text matching tasks and an impressive 19.75% increase in text-to-image matching tasks. These results underscore the cutting-edge and efficacious nature of our approach, highlighting its substantial advancements over existing methodologies. We also explore the auxiliary effect of using eye-gaze data of different scales on the model, finding that even a small portion of eye-gaze data can enhance the model’s multi-modal processing capability. Moreover, our framework performs feature alignment after the encoders, enhancing flexibility while reducing reliance on annotated data, laying a foundation for the future application of eye-tracking data in the medical domain. The code of this work will be available on Github [^1] https://github.com/Momarky/EGMA.
@@
  • EGMA outperforms existing state-of-the-art medical multi-modal pre-training methods, and realizes notable enhancements in image-text retrieval tasks.
  • EGMA demonstrates that even a small amount of eye-gaze data can effectively assist in multi-modal pre-training.
@@ III Method
Figure 2: The framework of EGMA. Part $A$ is the feature extraction module, aimed at obtaining a fine-grained similarity matrix for instances. Parts $B$ involves two types of guidance information, Gaze-guided Similarity (GS) and Gaze-guided Label (GL), derived from eye-gaze data. In parts $C$ and $D$, we propose two fine-grained multi-modal feature alignment methods guided by eye-gaze, leveraging the auxiliary information generated in part $B$ to further optimize the similarity matrix generated in part $A$.
As shown in Fig. 2, the framework of our proposed method consists of four main components. Firstly, we extract features from image and text in part $A$ to obtain a refined instance-level similarity matrix. Secondly, in part $B$, we integrate textual transcripts derived from radiologists’ audio, images, and eye-gaze data, to visualize and map radiologists’ attention onto specific regions of images during diagnosis. This process establishes alignment between texts and images, facilitating model training. The detailed gaze data processing methods are described in Sec. III-A. Given that eye-gaze data tightly links textual and localized visual information, after obtaining auxiliary information from part $B$, we introduce eye-gaze guided refined alignment training strategies, as depicted in Parts $C$ and $D$ of Fig. 2. Specifically, we introduce the optimization algorithm for eye-gaze guided fine-grained text-image similarity matrix in Part $C$ in Sec. III-B. Finally, in Sec. III-C, we present the algorithm for eye-gaze guided cross-modal mapping.
@@ III-A Multi-modal Data Processing
With the development of data collection technologies such as eye-tracking and speech recognition, it has become possible to collect and process multi-modal interaction data of radiologists during the diagnostic process. As illustrated in Fig. 3, the presentation of multi-modal data in the MIMIC-EYE [14] dataset includes radiologists’ audio, text transcript, eye-gaze data, and image. Since each modality is synchronized, the audio data is aligned with the eye-gaze data in time. By segmenting the audio based on the time before and after the pronunciation of each word, we can align the transcripts with the audio, thereby aligning word-level text with eye-gaze data. Subsequently, we generate attention heatmap based on eye-gaze data and images to represent the image regions the radiologist focuses on. Through the aforementioned data processing steps, we achieve precise alignment between word-level text and image regions. It is noteworthy that due to the rapid speech rate of radiologists, there may be no available eye-gaze data within the time interval corresponding to a single word. In Fig. 3, the word “with” in the transcript has no corresponding gaze data. Another common and unavoidable issue is the loss of eye-gaze data caused by blinking and intense head movement of radiologist, as seen in the last two words of the transcript. Due to these technical challenges, achieving perfect pairing between words and image regions is difficult. However, as shown in the right side of Fig. 3, adjusting the text to the sentence level largely mitigates the issue of missing word-level heatmap (Heatmap with red edge), and the semantic information of the entire sentence also encompasses the information of each word. Therefore, in this work, we process text features at the sentence level.
@@ III-B Eye-gaze Guided Fine-grained Alignment
The core idea of contrastive learning is to bring the features of related samples closer while pushing away the features of unrelated samples. During the training progress of CLIP model, assuming a batch size of $b$ and input data $\left\{x_{k}^{I},x_{k}^{T}\right\}$ ($k=1,\cdots,b$) representing image-text pairs, global features $z_{k}^{I}=E_{I}(x_{k}^{I})\in\mathbb{R}^{1\times d}$ and $z_{k}^{T}=E_{T}(x_{k}^{T})\in\mathbb{R}^{1\times d}$ are obtained through image encoder $E_{I}$ and text encoder $E_{T}$. Subsequently, the cosine similarity $s_{k,l}^{I2T}$ and $s_{k,l}^{T2I}$ between the two modalities is computed, with the following formula:
@@
where $s_{k,l}^{I2T}$ is the image-to-text similarity, $s_{k,l}^{T2I}$ is the text-to-image similarity, and $l$ is the index number of the another modality. Then, the image-to-text contrastive loss $L_{k}^{I2T}$ for $x_{k}^{I}$ can be formulated as:
@@
where $\tau$ is learned temperature. Similarly, the text-to-image contrastive loss $L_{k}^{T2I}$ for $x_{k}^{T}$ is:
$L_{k}^{T2I}(x_{k}^{T},\left\{x_{l}^{I}\right\}_{l=1}^{b})=-\frac{1}{b}log\frac{exp(s_{k,k}^{T2I}/\tau)}{\textstyle\sum_{l}(exp(s_{k,l}^{T2I}/\tau))}$ (3)
It is worth noting that in the calculation of the loss mentioned above, both the image and text utilize global-level features, while the auxiliary information generated from eye-gaze data emphasizes the local-level features between modalities. Therefore, based on [4], we replace instance feature $z_{k}^{I}$ and $z_{k}^{T}$ with $P_{k}^{n}\in\mathbb{R}^{n\times d}$ and $S_{k}^{m}\in\mathbb{R}^{m\times d}$, where $P_{k}^{i}$($1\leqslant i\leqslant n$) is the $i$-th patch feature of $x_{k}^{I}$ and $S_{k}^{j}$($1\leqslant j\leqslant m$) is the $j$-th sentence feature of $x_{k}^{T}$, and $n,m$ are the image patch number and the sentence number of report. Then we calculate the similarities of sentence-to-patch $x_{k}^{S2P}\in\mathbb{R}^{m\times n}$ and patch-to-sentence $x_{k}^{P2S}\in\mathbb{R}^{n\times m}$ in one instance:
@@
$fL_{k}^{S2P}=mlce(x_{k}^{S2P},GL_{k}),\ fL_{k}^{P2S}=mlce(x_{k}^{P2S},(GL_{k})^{\rm T})$ (5)
where $mlce$ is the multi-label cross-entropy loss. Subsequently, we calculate the fine-grained features $\hat{z}_{k}^{I}=Mean_{i}(Max_{j}(x_{k}^{P2S}))$ and $\hat{z}_{k}^{T}=Mean_{j}(Max_{i}(x_{k}^{S2P}))$. Then, we replace the $z_{k}^{I},z_{k}^{T}$ with the updated $\hat{z}_{k}^{I},\hat{z}_{k}^{T}$ in Eq. 1. Finally, the fine-grained image-to-text loss $\hat{L}_{k}^{I2T}$ and text-to-image loss $\hat{L}_{k}^{T2I}$ are computed based on Eq. 2 and Eq. 3. The formula for our Eye-gaze Guided Fine-grained (EGF) alignment loss is as follows:
@@
III-C Eye-gaze Guided Cross-model Mapping
In the previous section, we replaced the global instance logits in the traditional batch clip loss with fine-grained instance logits that consider local features and optimized the alignment between these local features using gaze information. The text in our work is recorded by radiologists while observing images, implying a close semantic relationship between the focus region and the corresponding text. To further optimize the alignment between modalities, we continue to incorporate eye-gaze data assistance into the cross-modal mapping process. In this work, we first utilize matrices $GS_{k}$ and $x_{k}^{P2S}$ to generate the image-to-text alignment weight matrix $W^{I2T}\in\mathbb{R}^{n\times m}$. The calculation formula is as follows:
@@ III-C Eye-gaze Guided Cross-model Mapping
where $norm$ is normalization and $\omega$ consists of sparse and binarize operations. Similarly, the text-to-image alignment weight matrix $W^{T2I}\in\mathbb{R}^{m\times n}$ is:
$W^{T2I}=norm(\omega(x_{k}^{S2P})+(GS_{k})^{\rm T})$ (8)
After obtaining the weight matrix, we perform the mapping from text features $S_{k}^{m}$ to image features $Cross\_P_{k}^{n}\in\mathbb{R}^{n\times d}$ and from image features $P_{k}^{n}$ to text features $Cross\_S_{k}^{m}\in\mathbb{R}^{m\times d}$ according to the following formula:
@@
Finally, the total loss of our model within a batch is $L=L_{EGF}+L_{EGM}$. In our training process, considering the proportion of eye-gaze data, batches may contain both types of data. When encountering samples without eye-gaze data, the EGF module does not compute the loss from Eq. 5, and the weight matrix in the Eq. 7 and Eq. 8 of EGM module also excludes the $GS_{k}$.

### v3

# Eye-gaze Guided Multi-modal Alignment for Medical Representation Learning

@@ Abstract
In the medical multi-modal frameworks, the alignment of cross-modality features presents a significant challenge. However, existing works have learned features that are implicitly aligned from the data, without considering the explicit relationships in the medical context. This data-reliance may lead to low generalization of the learned alignment relationships. In this work, we propose the Eye-gaze Guided Multi-modal Alignment (EGMA) framework to harness eye-gaze data for better alignment of medical visual and textual features. We explore the natural auxiliary role of radiologists’ eye-gaze data in aligning medical images and text, and introduce a novel approach by using eye-gaze data, collected synchronously by radiologists during diagnostic evaluations. We conduct downstream tasks of image classification and image-text retrieval on four medical datasets, where EGMA achieved state-of-the-art performance and stronger generalization across different datasets. Additionally, we explore the impact of varying amounts of eye-gaze data on model performance, highlighting the feasibility and utility of integrating this auxiliary data into multi-modal alignment framework.
@@ I Introduction
With the development of multi-modal learning, pre-trained models can now utilize large amounts of paired multi-modal data, such as image-text pairs, audio-text pairs, etc., to optimize the multi-modal feature extraction and alignment capabilities. With the emergence of the CLIP [1] model, contrastive learning has become the prominent framework of multi-modal learning. The advantage of this framework lies in its simplicity of structure and it does not require sample-level annotations. However, the main drawback is its heavy reliance on the scale of training data. Subsequent works have optimized this framework by leveraging potential auxiliary information between image and text data. For instance, GLIP [2] and RegionCLIP [3] utilized pre-predicted annotation information to perform fine-grained region-level pre-training. They introduced detection networks firstly to predict image regions relevant to the text prompt, and then trained the model to align these image regions with their corresponding text descriptions. However, these models heavily rely on the performance of the ROI detector and have high computational complexity. FILIP [4] proposed a refined multi-modal alignment operation after the encoder, relying solely on image patches and text tokens. Although this further explores the local feature relationships between multi-modal data, it still requires sufficient data support. When training on small-scale datasets, especially in the medical field, accurately learning alignment features between modalities becomes more challenging [5, 6].
To address the scarcity of medical data, studies [7, 8] have introduced self-supervised training into the CLIP framework to further enhance encoder performance. Additionally, weak labels between images and texts have been incorporated during pre-training to aid multi-modal alignment [9]. Some studies [10, 11] utilized fine-grained alignment between chest image patches and text tokens for pre-training [4]. However, unlike natural images and text, the relationship between medical images and diagnostic text is often more complex and challenging to learn. Moreover, with insufficient data, models are prone to learning shortcut features unrelated to disease diagnosis, resulting in poor generalization ability [12, 13, 14]. Therefore, it is crucial to learn useful alignment information from relatively limited medical multi-model datasets.
@@
In this study, we fully explore the auxiliary role of eye-gaze data from radiologists in multi-model alignment. Eye-gaze data can intuitively reflect the image regions radiologists focus on, providing insights into their cognitive behavior during diagnosis [15]. Therefore, compared to refined annotations like bounding boxes and masks, eye-gaze data can also provide useful auxiliary information for the model [16, 17, 14]. Moreover, collecting eye-gaze data from radiologists during the diagnostic process is more time-efficient than annotating bounding boxes and masks [18, 14]. For the multi-modal medical dataset, EYE GAZE [19] and REFLACX [20] collected eye-gaze data from radiologists while diagnosing chest X-rays. Additionally, these datasets recorded synchronized voice data, where radiologists verbalized their diagnoses while observing the images. As shown in Fig. 1, we found that the radiologists’ attention regions on the image naturally align with the diagnostic text over time. Therefore, we believe this type of eye-gaze data can provide expert prior knowledge for training the alignment between medical visual and textual features. Thus, considering the utilization of eye-gaze data to assist in multi-modal model training, we propose the Eye-gaze Guided Multi-modal Alignment framework (EGMA). Our model first segments the transcribed text into individual sentences and obtains radiologists’ attention heatmaps. Subsequently, we obtain encoded features of image patches and sentences through image and text encoders, generating instance-level similarity matrix. Then, we compute the loss between this matrix and the attention heatmaps, integrating refined feature representations for subsequent contrastive loss. To further leverage the assisting role of eye-gaze data in aligning images and texts, we combine the eye-gaze heatmaps with the similarity matrix derived from model, serving as weights to calculate cross-modality mapping loss. Experimental results on zero-shot classification and retrieval tasks reveal that our framework surpasses other leading methods in performance across diverse datasets and under multiple dataset size scenarios. Specifically, the EGMA framework yielded a remarkable 3.9% improvement in image-to-text matching tasks and an impressive 19.75% increase in text-to-image matching tasks. These results underscore the cutting-edge and efficacious nature of our approach, highlighting its substantial advancements over existing methodologies. We also explore the auxiliary effect of using eye-gaze data of different scales on the model, finding that even a small portion of eye-gaze data can enhance the model’s multi-modal processing capability. Moreover, the fine-tuned classification results of EGMA achieved the best performance across multiple datasets. The code of this work is available on  [^1] https://github.com/Momarky/EGMA.
@@
  • EGMA outperforms existing state-of-the-art medical multi-modal pre-training methods, and realizes notable enhancements in image classification and image-text retrieval tasks.
  • EGMA demonstrates that even a small amount of eye-gaze data can effectively assist in multi-modal pre-training and improve the feature representation ability of the model.
@@ III Method
Figure 2: The framework of EGMA. After images and text are processed by the encoder in Part $A$, patch feature and sentence feature representations are obtained, resulting in a fine-grained similarity matrix for instances. Subsequently, the two types of eye-gaze-based auxiliary information obtained in Part $B$ are used for fine-grained and cross-mapping alignment in Part $C$ and Part $D$, respectively.
As shown in Fig. 2, the framework of our proposed method consists of four main components. Firstly, we extract features from image and text in part $A$ to obtain a refined instance-level similarity matrix. Secondly, in part $B$, we integrate textual transcripts derived from radiologists’ audio, images, and eye-gaze data, to visualize and map radiologists’ attention onto specific regions of images during diagnosis. This process establishes alignment between texts and images, facilitating model training. The detailed gaze data processing methods are described in Sec. III-A. Given that eye-gaze data tightly links textual and localized visual information, after obtaining auxiliary information from part $B$, we introduce eye-gaze guided refined alignment training strategies, as depicted in Parts $C$ and $D$ of Fig. 2. Specifically, we introduce the optimization algorithm for eye-gaze guided fine-grained text-image similarity matrix in Part $C$ in Sec. III-B. Finally, in Sec. III-C, we present the algorithm for eye-gaze guided cross-modality mapping.
@@ III-A Multi-modal Data Processing
With the development of data collection technologies such as eye-tracking and speech recognition, it has become possible to collect and process multi-modal interaction data of radiologists during the diagnostic process. In this work, we utilize MIMIC-EYE [33] datasets as our training set, consisting of 3689 images extracted from the MIMIC datasets [34, 35, 36, 37]. Each sample is accompanied by corresponding eye-tracking data and transcripts text. These eye-tracking data are provided by the publicly available EYE GAZE [19] and REFLACX [20] datasets on PhysioNet [38]. Since each modality is synchronized, the audio data is aligned with the eye-gaze data in time. By segmenting the audio based on the time before and after the pronunciation of each word, we can align the transcripts with the audio, thereby aligning sentence-level text with eye-gaze data. Subsequently, we generate attention heatmap based on eye-gaze data and images to represent the image regions the radiologist focuses on. Through the aforementioned data processing steps, we achieve precise alignment between sentence-level text and image regions. Detailed processing method of eye-gaze and audio transcripts can be found at Supplementary Materials.
@@ III-B Eye-gaze Guided Fine-grained Alignment
The core idea of contrastive learning is to bring the features of related samples closer while pushing away the features of unrelated samples. During the training progress of CLIP [1] model, assuming a batch size of $b$ and input data $\left\{x_{k}^{I},x_{k}^{T}\right\}$ ($k=1,\cdots,b$) representing image-text pairs, global features $z_{k}^{I}=E_{I}(x_{k}^{I})\in\mathbb{R}^{1\times d}$ and $z_{k}^{T}=E_{T}(x_{k}^{T})\in\mathbb{R}^{1\times d}$ are obtained through image encoder $E_{I}$ and text encoder $E_{T}$. Subsequently, the cosine similarity $s_{k,l}^{I2T}$ and $s_{k,l}^{T2I}$ between the two modalities is computed, with the following formula:
@@
where $s_{k,l}^{I2T}$ is the image-to-text similarity, $s_{k,l}^{T2I}$ is the text-to-image similarity, and $l$ is the index number of the another modality. Then, the image-to-text contrastive loss $L_{k}^{I2T}$ for $x_{k}^{I}$ and text-to-image contrastive loss $L_{k}^{T2I}$ for $x_{k}^{T}$ can be formulated as:
@@
$L_{k}^{T2I}(x_{k}^{T},\left\{x_{l}^{I}\right\}_{l=1}^{b})=-\frac{1}{b}log\frac{exp(s_{k,k}^{T2I}/\tau)}{\textstyle\sum_{l}(exp(s_{k,l}^{T2I}/\tau))}$
where $\tau$ is a learned temperature. It is worth noting that in the calculation of the loss mentioned above, both the image and text utilize global-level features, while the auxiliary information generated from eye-gaze data emphasizes the local-level features between modalities. Therefore, based on [4], we replace instance feature $z_{k}^{I}$ and $z_{k}^{T}$ with $P_{k}^{n}\in\mathbb{R}^{n\times d}$ and $S_{k}^{m}\in\mathbb{R}^{m\times d}$, where $P_{k}^{i}$($1\leqslant i\leqslant n$) is the $i$-th patch feature of $x_{k}^{I}$ and $S_{k}^{j}$($1\leqslant j\leqslant m$) is the $j$-th sentence feature of $x_{k}^{T}$, and $n,m$ are the image patch number and the sentence number of report. Then we calculate the similarities of sentence-to-patch $x_{k}^{S2P}\in\mathbb{R}^{m\times n}$ and patch-to-sentence $x_{k}^{P2S}\in\mathbb{R}^{n\times m}$ in one instance:
@@
$fL_{k}^{S2P}=mlce(x_{k}^{S2P},GL_{k})$ (4)
$fL_{k}^{P2S}=mlce(x_{k}^{P2S},(GL_{k})^{\rm T})$
where $mlce$ is the multi-label cross-entropy loss. Subsequently, we calculate the fine-grained features $\hat{z}_{k}^{I}=Mean_{i}(Max_{j}(x_{k}^{P2S}))$ and $\hat{z}_{k}^{T}=Mean_{j}(Max_{i}(x_{k}^{S2P}))$. Then, we replace the $z_{k}^{I},z_{k}^{T}$ with the updated $\hat{z}_{k}^{I},\hat{z}_{k}^{T}$ in Eq. 1. Finally, the fine-grained image-to-text loss $\hat{L}_{k}^{I2T}$ and text-to-image loss $\hat{L}_{k}^{T2I}$ are computed based on Eq. 2. The formula for our Eye-gaze Guided Fine-grained (EGF) alignment loss is as follows:
@@
III-C Eye-gaze Guided Cross-modality Mapping
In the previous section, we replaced the global instance logits in the traditional batch clip loss with fine-grained instance logits that consider local features and optimized the alignment between these local features using gaze information. The text in our work is recorded by radiologists while observing images, implying a close semantic relationship between the focus region and the corresponding text. To further optimize the alignment between modalities, we continue to incorporate eye-gaze data assistance into the cross-modality mapping process. In this work, we first utilize matrices $GS_{k}$, $x_{k}^{P2S}$ and $x_{k}^{S2P}$ to generate the image-to-text and text-to-image alignment weight matrix $W^{I2T}\in\mathbb{R}^{n\times m}$ and $W^{T2I}\in\mathbb{R}^{m\times n}$. The calculation formula is as follows:
@@ III-C Eye-gaze Guided Cross-modality Mapping
$W^{T2I}=norm(\omega(x_{k}^{S2P})+(GS_{k})^{\rm T})$
where $norm$ is normalization and $\omega$ consists of sparse and binarize operations. After obtaining the weight matrix, we perform the mapping from text features $S_{k}^{m}$ to image features $Cross\_P_{k}^{n}\in\mathbb{R}^{n\times d}$ and from image features $P_{k}^{n}$ to text features $Cross\_S_{k}^{m}\in\mathbb{R}^{m\times d}$ according to the following formula:
@@
Finally, the total loss of our model within a batch is $L=L_{EGF}+L_{EGM}$. In our training process, considering the proportion of eye-gaze data, batches may contain both types of data. When encountering samples without eye-gaze data, the EGF module does not compute the loss from Eq. 4, and the weight matrix in the Eq. 6 of EGM module also excludes the $GS_{k}$.